# Univariate-Guided Interaction Modeling

### v1

@@ 8 Discussion
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[|\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\big(|\widehat{\beta}_{0,k}-\widehat{\beta}_{0,k}^{(-i)}|+|\widehat{\beta}_{1,k}-\widehat{\beta}_{1,k}^{(-i)}||X_{ik}|\big)\Big]$
@@
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\Biggl\{\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[|\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\big(|\widehat{\beta}_{0,k}-\widehat{\beta}_{0,k}^{(-i)}|+|\widehat{\beta}_{1,k}-\widehat{\beta}_{1,k}^{(-i)}||X_{ik}|\big)\Big]$
@@
$|\widehat{\theta}_{j}-\theta_{j}^{*}|$ $\leq\frac{|\widehat{\beta}_{j}-\beta_{j}^{*}|}{|\widehat{\beta}_{1,j}|}+|\beta_{j}^{*}||\frac{1}{\beta_{1,j}^{*,\mathrm{uni}}}-\frac{1}{\widehat{\beta}_{1,j}}|$ (12)
@@
n v And if $2D_{1}\leq U_{1}$, then
$|\widehat{\theta}_{j}-\theta_{j}^{*}|$ $\leq\frac{2\Delta}{|\beta_{1,j}^{*,\mathrm{uni}}|}+|\beta_{j}^{*}|\frac{2D_{j,1}}{(\beta_{1,j}^{*,\mathrm{uni}})^{2}}$ (14)
@@
$\sum_{j\in S_{M}}R_{j}$ $\geq-\frac{O_{a}(1)\Delta}{U_{1}}\Biggl\{(D_{0}+D_{1})Q_{1}+(D_{0}+Q_{3}M_{1})(\frac{\Delta D_{0}}{U_{1}}+\frac{D_{0}D_{1}}{U_{1}^{2}}+D_{0})$ (15)
$+(D_{0}Q_{3}+M_{1}Q_{2})(\frac{\Delta D_{1}}{U_{1}}+\frac{D_{1}^{2}}{U_{1}^{2}}+D_{1})+D_{1}\Delta Q_{2}+\Delta D_{0}+\Delta D_{1}Q_{3}+D_{0}\Delta Q_{3}\Biggr\}$ (16)
@@
with $\widetilde{D}=\max(D_{0},D_{1})$ and $\widetilde{Q}=\max(M_{1},Q_{1},Q_{2},Q_{3})$. Let $A_{3}=(\widetilde{D}+\widetilde{D}^{2}+\widetilde{D}^{3})(\widetilde{Q}+\widetilde{Q}^{2})$. Then
@@
$\mathbb{P}(\tilde{Q}>A)$ $\leq O_{a}(1)(n+1)|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A^{2})+O_{a}(1)(|S_{M}|+1)\exp(-\Omega_{a}(1)nA^{2})$ (18)
$+O_{a}(1)(|S_{M}|+1)\exp(-\Omega_{a}(1)n^{3/4}A)$ (19)
Hence,
$\mathbb{P}(\tilde{Q}>A)$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)$ (21)
Recall that $\widetilde{Q_{1}}=\max_{k\in S_{M}^{\prime}}|P_{n}[\varepsilon_{i}^{\prime}X_{ik}]|$ and $\delta=\max_{k\in S_{M}^{\prime}}|\mathbb{E}(\epsilon_{i}^{\prime}X_{ik})|$. So,
@@
$\mathbb{P}(A_{3}>(t+t^{2}+t^{3})(A+A^{2}))$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$ (23)
So for all $t\in(0,O_{a}(1))$ and $A=\Omega_{a}(1)$, we have
$\mathbb{P}(A_{3}>tA^{2})$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$ (24)
Therefore, for all $t\in(0,O_{a}(1))$, we have
@@
Therefore, we conclude that for all $t\in(0,O_{a}(1))$,
$\mathbb{P}\Big(E^{c}\cap\big(\Delta>O_{a}(1)\frac{\lambda_{1}+\delta+t}{\eta^{*}_{M}-O_{a}(1)n^{-\epsilon/2}-O_{a}(1)t}\big)\Big)$ (25)
$\leq O_{a}(1)|S_{M}|\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}\right)+O_{a}(1)|S_{M}|\exp\!\left(-\Omega_{a}(1)n^{3/4}\right)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$ (26)
$+O_{a}(1)\exp(-\Omega_{a}(1)n^{1-\epsilon})$ (27)
@@
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates adapted according to the bounds established in 1, 3, 4 and 5. In particular, the following rates are modified :
@@
  • Let $A_{3}^{\prime}=(\widetilde{D^{\prime}}+\widetilde{D^{\prime}}^{2}+\widetilde{D^{\prime}}^{3})(\widetilde{Q^{\prime}}+\widetilde{Q^{\prime}}^{2})$. Then for all $t\in(0,O_{a}(1))$, we have
    $\mathbb{P}(A_{3}>t)\leq O_{a}(1)|S_{A}|n\exp({-\Omega_{a}(1)n^{\frac{3}{5}}t^{2}})$

### v2

@@ 8 Discussion
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[|\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\big(|\widehat{\beta}_{0,k}^{\mathrm{uni}}-\widehat{\beta}_{0,k}^{(-i)}|+|\widehat{\beta}_{1,k}^{\mathrm{uni}}-\widehat{\beta}_{1,k}^{(-i)}||X_{ik}|\big)\Big]$
@@
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\Biggl\{\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[|\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\big(|\widehat{\beta}_{0,k}^{\mathrm{uni}}-\widehat{\beta}_{0,k}^{(-i)}|+|\widehat{\beta}_{1,k}^{\mathrm{uni}}-\widehat{\beta}_{1,k}^{(-i)}||X_{ik}|\big)\Big]$
@@
$|\widehat{\theta}_{j}-\theta_{j}^{*}|$ $\leq\frac{|\widehat{\beta}_{j}-\beta_{j}^{*}|}{|\widehat{\beta}_{1,j}|}+|\beta_{j}^{*}|\big|\frac{1}{\beta_{1,j}^{*,\mathrm{uni}}}-\frac{1}{\widehat{\beta}_{1,j}}\big|$
@@
And if $2D_{1}\leq U_{1}$, then
$|\widehat{\theta}_{j}-\theta_{j}^{*}|$ $\leq\frac{2\Delta}{|\beta_{1,j}^{*,\mathrm{uni}}|}+|\beta_{j}^{*}|\frac{2D_{j,1}}{(\beta_{1,j}^{*,\mathrm{uni}})^{2}}$
@@
$\sum_{j\in S_{M}}R_{j}$ $\geq-\frac{O_{a}(1)\Delta}{U_{1}}\Biggl\{(D_{0}+D_{1})Q_{1}+(D_{0}+Q_{3}M_{1})(\frac{\Delta D_{0}}{U_{1}}+\frac{D_{0}D_{1}}{U_{1}^{2}}+D_{0})$
$+(D_{0}Q_{3}+M_{1}Q_{2})(\frac{\Delta D_{1}}{U_{1}}+\frac{D_{1}^{2}}{U_{1}^{2}}+D_{1})+D_{1}\Delta Q_{2}+\Delta D_{0}+\Delta D_{1}Q_{3}+D_{0}\Delta Q_{3}\Biggr\}$
@@
with $\widetilde{D}=\max(D_{0},D_{1})$ and $\widetilde{Q}=\max(M_{1},Q_{1},Q_{2},Q_{3})$. Let $A_{3}=(\widetilde{D}+\widetilde{D}^{2}+\widetilde{D}^{3})(\widetilde{Q}+\widetilde{Q}^{2})$. Then,
@@
$\mathbb{P}(\tilde{Q}>A)$ $\leq O_{a}(1)(n+1)|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A^{2})+O_{a}(1)(|S_{M}|+1)\exp(-\Omega_{a}(1)nA^{2})$
$+O_{a}(1)(|S_{M}|+1)\exp(-\Omega_{a}(1)n^{3/4}A)$
Hence, $\mathbb{P}(\tilde{Q}>A)\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)$. Recall that $\widetilde{Q_{1}}=\max_{k\in S_{M}^{\prime}}|P_{n}[\varepsilon_{i}^{\prime}X_{ik}]|$ and $\delta=\max_{k\in S_{M}^{\prime}}|\mathbb{E}(\epsilon_{i}^{\prime}X_{ik})|$. So,
@@
$\mathbb{P}(A_{3}>(t+t^{2}+t^{3})(A+A^{2}))$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
So for all $t\in(0,\Omega_{a}(1))$ and $A=\Omega_{a}(1)$, we have
$\mathbb{P}(A_{3}>tA^{2})$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
Therefore, for all $t\in(0,\Omega_{a}(1))$, we have
@@
Therefore, we conclude that for all $t\in(0,\Omega_{a}(1))$,
$\mathbb{P}\Big(E^{c}\cap\big(\Delta>O_{a}(1)\frac{\lambda_{1}+\delta+t}{\eta^{*}_{M}-O_{a}(1)n^{-\epsilon/2}-O_{a}(1)t}\big)\Big)$
$\leq O_{a}(1)|S_{M}|\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}\right)+O_{a}(1)|S_{M}|\exp\!\left(-\Omega_{a}(1)n^{3/4}\right)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
$+O_{a}(1)\exp(-\Omega_{a}(1)n^{1-\epsilon})$
@@
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates adapted according to the bounds established in Lemmas 1, 3, 4 and 5. In particular, the following rates are modified :
@@
  • Let $A_{3}^{\prime}=(\widetilde{D^{\prime}}+\widetilde{D^{\prime}}^{2}+\widetilde{D^{\prime}}^{3})(\widetilde{Q^{\prime}}+\widetilde{Q^{\prime}}^{2})$. Then for all $t\in(0,\Omega_{a}(1))$, we have
    $\mathbb{P}(A_{3}>t)\leq O_{a}(1)|S_{A}|n\exp({-\Omega_{a}(1)n^{3/5}t^{2}})$